OpenStudy (anonymous):

v(t)=90-30ln(3t+1) solve for t

OpenStudy (amistre64):

remember the "e" undoes the "ln"

OpenStudy (anonymous):

so far i got 3=ln(3t+1)

OpenStudy (anonymous):

now im stuck

OpenStudy (anonymous):

kk got it

OpenStudy (anonymous):

thanks :)

OpenStudy (amistre64):

"e" both sides e^3 = e^(ln(3t+1)) ; since e cancels ln e^3 = 3t + 1 and the rest is back to simpler stuff

OpenStudy (amistre64):

youre welcome :)